3a+2[2+3(9-1)]=2(3a+4)
Answer by Lal(16)

Parentheses
Exponents
Multiply
Divide
Add
Subtract

3a + 2 [2 + 3(9 - 1)] = 2(3a + 4)
3a + 2 [2 + 3 (8)] = 2 (3a + 4)
3a + 2 [2 + 24] = 6a + 8
3a + 2 [26] = 6a + 8
3a + 52 = 6a + 8
-3a = -44
a = 14 and 2/3